You have a box of chocolates that contains 51 pieces of which 29 are solid chocolate, 12 are filled with cashews, and 10 are filled with cherries. All the candies look exactly alike. You select a piece, eat it, select a second piece, eat it, and finally eat one last piece. Find the probability of selecting a solid chocolate piece followed by two cherry-filled chocolates.
Answer by greenestamps(13198)



1st selection: 51 choices; 29 are solid chocolate. P(solid) = 29/51

2nd selection: 50 chocolates left; 10 are cherry. P(cherry) = 10/50

3rd selection: 49 chocolates left; 9 are cherry. P(cherry) = 9/49

P(solid then cherry then another cherry) = (29/51)(10/50)(9/49)

Perform the calculation and express the result in the required/desired form.